# Determining Energy Efficiency Sweet Spots in Production LLM Inference

### v1

@@ 4.1.1. Prefill Stage FLOPs
Input projections to $Q$, $K$, $V$: the input $X\in\mathbb{R}^{n_{\text{in}}\times d}$ is multiplied by three weight matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$:
$F_{\text{proj}}=3\cdot(2n_{\text{in}}d^{2})=6n_{\text{in}}d^{2}.$
@@
$F_{\text{att}}(n_{\text{in}})=8n_{\text{in}}d^{2}+4n_{\text{in}}^{2}d.$
@@
$F_{\text{FFN}}(n_{\text{in}})=8n_{\text{in}}d^{2}+8n_{\text{in}}d^{2}=16n_{\text{in}}d^{2}.$
corresponding to $XW_{1}$ and the subsequent $W_{2}$ projection back to dimension $d$.
@@
$F_{\text{prefil, layer}}(n_{\text{in}})=24n_{\text{in}}d^{2}+4n_{\text{in}}^{2}d.$
@@
$F_{\text{prefill}}=L\cdot\big(24n_{\text{in}}d^{2}+4n_{\text{in}}^{2}d\big).$
@@ 4.1.2. Decode Stage FLOPs
$F_{\text{decode, layer}}(t)=24d^{2}+4(n_{\text{in}}+t-1)d.$
@@ 4.1.3. Total Inference FLOPs
$F_{\text{total}}(n_{\text{in}},n_{\text{out}})$ $=F_{\text{prefill}}+F_{\text{decode}}$
@@
$F_{\text{total}}(n_{\text{in}},n_{\text{out}})=2\,L\,d\;\Big(12\,d\,n_{\text{in}}+12\,d\,n_{\text{out}}+2\,n_{\text{in}}^{2}+2\,n_{\text{out}}\,n_{\text{in}}+n_{\text{out}}^{2}-n_{\text{out}}\Big).$
We note that the external factors L and d depends only on model size (depth and width) and are constant for a given model. Since our goal is to capture how efficiency varies as a function of the input length $n_{\text{in}}$ and the output length $n_{\text{out}}$, we remove this constant multipliers and express the energy as a linear combination of polynomial terms in $n_{\text{in}}$ and $n_{\text{out}}$.
$E_{\text{tot}}(n_{\text{in}},n_{\text{out}})=\theta_{0}\,n_{\text{out}}+\theta_{1}\,n_{\text{in}}^{2}+\theta_{2}\,n_{\text{out}}\,n_{\text{in}}+\theta_{3}\,n_{\text{in}}+\theta_{4}\,n_{\text{out}}^{2}+\theta_{5}\,n_{\text{out}}.$
@@
To analyze the computational cost on a per-output-token basis, we divide the total energy by the number of generated tokens $n_{\text{out}}$. This provides the average energy required to produce a single token, which is useful for comparing efficiency across different input–output configurations. By aggregating $\theta_{5}$ to $\theta_{0}$, we obtain:
(4) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}\,n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}\,n_{\text{in}}+\frac{\theta_{3}\,n_{\text{in}}}{n_{\text{out}}}+\theta_{4}\,n_{\text{out}}.$
@@
4.2. Memory and FLOPs-based Model
While FLOPs capture the arithmetic workload of inference, modern accelerators increasingly operate in a regime where energy consumption is dominated not by computation but by memory traffic, in particular the movement of activations, weights, and KV-cache tensors across the memory hierarchy. To accurately model energy, we therefore incorporate the memory-access cost alongside the computational cost derived in the previous section. In the following subsections, we present an analytical decomposition of the memory contribution during both the prefill and decode phases. As before, we analyze a single Transformer layer with hidden dimension $d$ and input length $n_{\text{in}}$, and $n_{q}$ number of heads . We again omit the negligible contributions of bias reads, layer-normalization reads/writes, and softmax.
@@ 4.2.1. Prefill Stage Memory Operations
$M_{\text{proj}}=2n_{\text{in}}d+d^{2}.$
@@
$M_{\text{att}}(n_{\text{in}})=8n_{\text{in}}d+2d^{2}+2n_{\text{in}}^{2}n_{q}.$
@@
$M_{\text{prefill}}=L(M_{\text{att}}+M_{\text{FFN}})=L\left(10n_{\text{in}}d+10d^{2}+2n_{\text{in}}^{2}n_{q}\right).$
@@ 4.2.3. Total Memory Operations
$M_{\text{total}}(n_{\text{in}},n_{\text{out}})$ $=M_{\text{prefill}}+M_{\text{decode}}$
@@
As for the FLOPs-based model, we express the energy consumption as a linear combination of polynomial terms in $n_{\text{in}}$ and $n_{\text{out}}$:
$E_{\text{tot}}(n_{\text{in}},n_{\text{out}})=\theta_{0}\,n_{\text{out}}+\theta_{1}\,n_{\text{in}}^{2}+\theta_{2}\,n_{\text{out}}\,n_{\text{in}}+\theta_{3}\,n_{\text{in}}+\theta_{4}\,n_{\text{out}}^{2}+\theta_{5}\,n_{\text{out}}+\theta_{6}.$
Relative to the $E_{\text{tot}}$ of the FLOPs model, this formulation introduces an additional constant term, $\theta_{6}$. This term accounts for the contribution arising from the $10d^{2}$ component in the $M_{\text{prefill}}$ expression, which is independent of both $n_{\text{in}}$ and $n_{\text{out}}$. Analyzing the energy consumption per-output-token and aggregating $\theta_{5}$ to $\theta_{0}$, we obtain:
(5) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}\,n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}\,n_{\text{in}}+\frac{\theta_{3}\,n_{\text{in}}}{n_{\text{out}}}+\theta_{4}\,n_{\text{out}}+\frac{\theta_{5}}{n_{\text{out}}}.$
4.3. Energy Efficiency Sweet Spots
Using the computational model introduced in Equation 5, we identify energy-efficiency Sweet Spots by considering the per-token energy as a function of the number of generated tokens $n_{\text{out}}$. To this end, we compute the derivative of $E_{\text{tok}}$ with respect to $n_{\text{out}}$ and set it equal to zero:
$\frac{\partial E_{\text{tok}}}{\partial n_{\text{out}}}=-\frac{\theta_{1}n_{\text{in}}^{2}+\theta_{3}n_{\text{in}}+\theta_{5}}{n_{\text{out}}^{2}}+\theta_{4}.=0$
@@ 4.3. Energy Efficiency Sweet Spots
(a) $N_{\text{req}}=10$.
(b) $N_{\text{req}}=100$.
(c) $N_{\text{req}}=1000$.
Figure 1. Energy Efficiency $\frac{\text{Tokens}}{\text{Joule}}$ of Falcon 7B with different number of requests (10,100,1000).
Energy Efficiency of Falcon 7B with different number of requests (10,100,1000).
5. Experimental Methodology
This section details the experimental methodology adopted to validate the proposed analytical models and to characterize the energy consumption of LLMs on modern hardware. We first describe the monitoring framework employed for the data acquisition. Next, we outline the selection of LLMs, ensuring a diverse representation of sizes and attention mechanisms. Next, we define the quantitative metrics used to assess Energy Efficiency and validate our theoretical derivations. Finally, we present the configurations for the dataset generation, engine building process and benchmark execution.
5.1. Monitoring Framework
The monitoring framework is designed to collect and analyze GPU metrics during LLM execution. In particular, hardware metrics such as nvmlDeviceGetPowerUsage and NVML_CLOCK_GRAPHICS are sampled using the NVIDIA Management Library Python API (pynvml (Corporation, 2025)). These metrics are sampled at 500 ms intervals, published through the ExaMon plugin, and stored in the ExaMon database. For both data collection and transmission we employ an updated version of ExaMon (Bartolini et al., 2019), based on IoTdb Database (Wang et al., 2025).
5.2. LLM Selection
Table 2. LLM test set grouped into size-based families (1B–9B), with architectural details for each model.
@@ 5.2. LLM Selection
We consider a diverse set of LLMs, spanning from 1B to 9B parameters, as summarized in LABEL:tab:llm_set. In addition to model size, the selection includes architectures with different design choices and attention mechanisms—such as Multi-Head, Grouped-Query, and Multi-Query Attentions. To facilitate comparison, the models are grouped into three size-based families, ranging from XS models (1B–1.5B) to M models (6.7B–9B).
5.3. Metrics and Measurement
We quantify the energy consumption $E_{\text{tot}}$ as the discrete integration of the mean instantaneous GPU power, measured as described in subsection 5.1 over the measurement interval $\Delta t$, according to the following expression:
$E_{\text{tot}}=\sum_{i}{\left(\frac{P_{i}+P_{i+1}}{2}\right)\Delta t},$
where $P_{i}$ denotes the power usage recorded at each time step $i$. We empirically obtain the Energy-per-Token $E_{\text{tok}}$ by normalizing the energy consumption for the total amount of output tokens $n_{\text{out}}$:
$E_{\text{tok}}=\frac{E_{\text{tot}}}{n_{\text{out}}}\quad\left[\frac{\mathrm{\text{Joules}}}{\mathrm{\text{Token}}}\right].$
@@ 5.3. Metrics and Measurement
$E_{\text{eff}}=\frac{n_{\text{out}}}{E_{\text{tot}}}\quad\left[\frac{\mathrm{\text{Tokens}}}{\mathrm{\text{Joule}}}\right].$
This measure captures the productivity of the model with respect to its energy budget, facilitating clearer cross-model comparisons.
5.4. Dataset Generation
To benchmark TensorRT-LLM under controlled and reproducible conditions, we generated synthetic request datasets using the official utility script provided with TensorRT-LLM. For each dataset, token sequences were sampled using the Llama-3.1-8B-Instruct tokenizer, with input and output lengths drawn from fixed distributions defined by a specified mean and zero variance. This allowed us to precisely control the number of input and output tokens per request. We generated datasets for all combinations of input and output lengths ranging from 64 to 4096 tokens (in powers of two), and for different workload sizes with the number of requests set to 10, 100, and 1000.
5.5. Engines Configuration
@@
5.6. Benchmark Execution
Benchmark runs were executed with trtllm-bench in throughput mode, a dedicated TensorRT-LLM utility for evaluating LLM performance on a given dataset, designed to maximize sustained token generation rate under a fixed workload. Each run pairs a prebuilt engine with a corresponding synthetic dataset, ensuring that model configuration and workload characteristics remain constant throughout the measurement. We first assessed different number of requests (10, 100, and 1000). Subsequently, fixing the number of requests to 1000, we evaluated the LLMs listed in LABEL:tab:llm_set while varying the input and output sequence lengths from 64 to 4096 tokens. By reusing the same engine across datasets and varying only the dataset parameters, the reported results isolate the impact of input/output length and request count on inference throughput.
6. Experimental Results
6.1. Setup
@@
(a) XS sizes models.
(b) S sizes models.
(c) M sizes models.
Figure 2. Energy Efficiency of tested LLM considering $n_{\text{in}}=64$ and varying output tokens, per size family (XS, S, M).
@@
6.2. Number of Requests Selection
To assess the impact of the requests count $N_{\text{req}}$ (i.e. the number of queries used for evaluation) on Energy Efficiency, we considered multiple workload sizes. We excluded the single-request setting because our monitoring framework introduces a non-negligible timing error, preventing reliable measurements in that regime where execution times are shorter.
@@ 6.2. Number of Requests Selection
Finally, smaller workloads amplify the relative impact of measurement overhead and necessitate shorter sampling intervals, which can further amplify measurement noise and result in less reliable estimates. For these reasons, in the subsequent experiments we focus on the saturation scenario and fix $N_{\text{req}}$ at 1000. A more detailed analysis of how workload size interacts with batch-size is deferred to subsection 8.2.
6.3. Energy Efficiency Evaluation

### v2

@@ 4.1.1. Prefill Stage FLOPs
Input projections to Query, Key, Value ($Q$, $K$, $V$): the input $X\in\mathbb{R}^{n_{\text{in}}\times d}$ is multiplied by three weight matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$:
$F_{\text{proj}}=3\cdot(2n_{\text{in}}d^{2})=6n_{\text{in}}d^{2}$
@@
$F_{\text{att}}=8n_{\text{in}}d^{2}+4n_{\text{in}}^{2}d$
@@
$F_{\text{FFN}}=8n_{\text{in}}d^{2}+8n_{\text{in}}d^{2}=16n_{\text{in}}d^{2}$
corresponding to the attention output multiplication with $W_{1}$ and the subsequent $W_{2}$ projection back to dimension $d$.
@@
[…]
@@
$F_{\text{prefill}}=L\cdot\big(24n_{\text{in}}d^{2}+4n_{\text{in}}^{2}d\big)$
@@ 4.1.2. Decode Stage FLOPs
$F_{\text{decode, layer}}(t)$ $=F_{\text{proj+out}}+F_{QK^{T}+V}+F_{\text{FFN}}$
[…]
@@ 4.1.3. Total Inference FLOPs
$F_{\text{total}}=2\,L\,d\;\Big(12\,d\,n_{\text{in}}+12\,d\,n_{\text{out}}+2\,n_{\text{in}}^{2}+2\,n_{\text{out}}\,n_{\text{in}}+n_{\text{out}}^{2}-n_{\text{out}}\Big)$
We note that the external factors L and d depends only on model size (depth and width) and are constant for a given model. Since our goal is to capture how efficiency varies as a function of input length $n_{\text{in}}$ and output length $n_{\text{out}}$, we remove these constant multipliers and express the energy as a linear combination of polynomial terms in terms of $n_{\text{in}}$ and $n_{\text{out}}$. Aggregating similar terms, we obtain:
$E_{\text{tot}}(n_{\text{in}},n_{\text{out}})=\theta_{0}\,n_{\text{out}}+\theta_{1}\,n_{\text{in}}^{2}+\theta_{2}\,n_{\text{out}}\,n_{\text{in}}+\theta_{3}\,n_{\text{in}}+\theta_{4}\,n_{\text{out}}^{2}$
@@
To analyze the computational cost on a per-output-token basis, we divide the total energy by the number of generated tokens $n_{\text{out}}$. This provides the average energy required to produce a single token, which is useful for comparing efficiency across different input–output configurations.
(4) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}\,n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}\,n_{\text{in}}+\frac{\theta_{3}\,n_{\text{in}}}{n_{\text{out}}}+\theta_{4}\,n_{\text{out}}$
@@
4.2. SweetSpot Model
While FLOPs capture the arithmetic workload of inference, modern accelerators increasingly operate in a regime where energy consumption is dominated not by computation but by memory traffic, in particular, the movement of activations, weights, and KV-cache tensors across the memory hierarchy (Yuan et al., 2024). To accurately model energy, we therefore incorporate the memory-access cost alongside the computational cost derived in the previous section. In the following subsections, we present an analytical decomposition of the memory contribution during both the prefill and decode phases, denoted by $M$. As before, we analyze a single Transformer layer with hidden dimension $d$ and input length $n_{\text{in}}$, and $n_{q}$ number of heads. We again omit the negligible contributions of bias reads, layer-normalization reads/writes, and softmax.
@@ 4.2.1. Prefill Stage Memory Operations
$M_{\text{proj}}=2n_{\text{in}}d+d^{2}$
@@
$M_{\text{att}}=8n_{\text{in}}d+2d^{2}+2n_{\text{in}}^{2}n_{q}$
@@
For one layer:
$M_{\text{prefil, layer}}=M_{\text{att}}+M_{\text{FFN}}=10n_{\text{in}}d+10d^{2}+2n_{\text{in}}^{2}n_{q}$
@@
$M_{\text{prefill}}=L\left(10n_{\text{in}}d+10d^{2}+2n_{\text{in}}^{2}n_{q}\right)$
@@ 4.2.3. Total Memory Operations
$M_{\text{total}}$ $=M_{\text{prefill}}+M_{\text{decode}}$
@@
As for the FLOPs-only model, we express the energy consumption as a linear combination of polynomial terms in $n_{\text{in}}$ and $n_{\text{out}}$:
$E_{\text{tot}}(n_{\text{in}},n_{\text{out}})=\theta_{0}\,n_{\text{out}}+\theta_{1}\,n_{\text{in}}^{2}+\theta_{2}\,n_{\text{out}}\,n_{\text{in}}+\theta_{3}\,n_{\text{in}}+\theta_{4}\,n_{\text{out}}^{2}+\theta_{5}$
Relative to the $E_{\text{tot}}$ of the FLOPs model, this formulation introduces an additional constant term, $\theta_{6}$. This term accounts for the contribution arising from the $10d^{2}$ component in the $M_{\text{prefill}}$ expression, which is independent of both $n_{\text{in}}$ and $n_{\text{out}}$. Analyzing the energy consumption per-output-token, we obtain:
(5) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}\,n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}\,n_{\text{in}}+\frac{\theta_{3}\,n_{\text{in}}}{n_{\text{out}}}+\theta_{4}\,n_{\text{out}}+\frac{\theta_{5}}{n_{\text{out}}}$
4.3. Peak Energy Efficiency Prediction
Using the formulation of SweetSpot introduced in Equation 5, we identify peak energy-efficiency configurations by considering the per-token energy as a function of the number of generated tokens $n_{\text{out}}$. To this end, we compute the derivative of $E_{\text{tok}}$ with respect to $n_{\text{out}}$ and set it equal to zero:
$\frac{\partial E_{\text{tok}}}{\partial n_{\text{out}}}=-\frac{\theta_{1}n_{\text{in}}^{2}+\theta_{3}n_{\text{in}}+\theta_{5}}{n_{\text{out}}^{2}}+\theta_{4}=0$
@@ 4.3. Peak Energy Efficiency Prediction
5. Experimental Study
This section details the experimental study conducted to validate SweetSpot and to evaluate the energy consumption of LLMs inference on modern hardware. We describe the experimental design, including the monitoring framework, LLM selection, metrics, datasets, engine configurations, benchmarking procedure, and system setup. We then report and analyze the experimental results, evaluating the energy efficiency of the LLM test set and the accuracy of our proposed analytical model SweetSpot.
5.1. Experimental Design
This subsection formalizes the experimental protocol, describing the measurement pipeline and design choices adopted to ensure controlled and reproducible evaluation across all experiments.
5.1.1. Monitoring Framework
To collect and analyze GPU metrics during LLM inference, we employ an updated version of ExaMon (Bartolini et al., 2019), based on IoTdb Database (Wang et al., 2025). In particular, hardware metrics such as nvmlDeviceGetPowerUsage and NVML_CLOCK_GRAPHICS are sampled using the NVIDIA Management Library Python API (pynvml (Corporation, 2025)). These metrics are sampled at 500 ms intervals, published through the ExaMon plugin, and stored in the ExaMon database.
5.1.2. LLM Selection
Table 2. LLM test set grouped into size-based families (1B–9B parameters), with architectural details for each model.
@@ 5.1.2. LLM Selection
We consider a diverse set of LLMs, ranging from 1B to 9B parameters, as summarized in LABEL:tab:llm_set. Beyond model size, the selection includes architectures with different design choices and attention mechanisms, including Multi-Head Attention (MHA) and Grouped-Query Attention (GQA). For GQA, we also account for different KV grouping ratios e.g. GQA (4×) denotes four query heads sharing a single KV head. To facilitate comparison, the models are grouped into three size-based families, ranging from XS models (1B–1.5B parameters) to M models (6.7B–9B parameters).
5.1.3. Metrics and Measurement
We quantify the energy consumption $E_{\text{tot}}$ as the discrete integration of the mean instantaneous GPU power, measured as described in Section 5.1.1 over the measurement interval $\Delta t$, according to the following expression:
$E_{\text{tot}}=\sum_{i}{\left(\frac{P_{i}+P_{i+1}}{2}\right)\Delta t}$
where $P_{i}$ denotes the power usage recorded at each time step $i$. We empirically obtain the Energy-per-Token ($E_{\text{tok}}$) by normalizing the energy consumption by the total amount of output tokens $n_{\text{out}}$:
$E_{\text{tok}}=\frac{E_{\text{tot}}}{n_{\text{out}}}\quad\left[\frac{\mathrm{\text{Joules}}}{\mathrm{\text{Token}}}\right]$
@@ 5.1.3. Metrics and Measurement
$E_{\text{eff}}=\frac{n_{\text{out}}}{E_{\text{tot}}}\quad\left[\frac{\mathrm{\text{Tokens}}}{\mathrm{\text{Joule}}}\right]$
This measure captures the productivity of the model with respect to its energy budget, facilitating comparisons between models.
5.1.4. Dataset Generation
To benchmark TensorRT-LLM under controlled and reproducible conditions, we generated synthetic request datasets using the official utility script provided with TensorRT-LLM. For each dataset, token sequences were sampled using the Llama-3.1-8B-Instruct tokenizer, with input and output lengths drawn from fixed distributions defined by a specified mean and zero variance. This allowed us to precisely control the number of input and output tokens per request. We generated the dataset. [^2] The dataset is publicly available on Zenodo: doi.org/10.5281/zenodo.18714476 for all combinations of input and output lengths ranging from 64 to 4096 tokens (in powers of two), and for different workload sizes with the number of requests set to 10, 100, and 1000.
5.1.5. Engines Configuration
@@
5.1.6. Benchmark Execution
Benchmark runs were executed with trtllm-bench in throughput mode, a dedicated TensorRT-LLM utility for evaluating LLM performance on a given dataset, designed to maximize sustained token generation rate under a fixed workload. Each run pairs a prebuilt engine with a corresponding synthetic dataset, ensuring that model configuration and workload characteristics remain constant throughout the measurement. We first assessed different number of requests (10, 100, and 1000) using Falcon-RW 7B. Subsequently, we evaluated the LLMs listed in LABEL:tab:llm_set by fixing the number of requests to 1000, while varying the input and output sequence lengths from 64 to 4096 tokens. By reusing the same engine across datasets and varying only the dataset parameters, the reported results isolate the impact of input/output length and request count on inference throughput.
5.1.7. System Setup
@@
(a) $N_{\text{req}}=10$.
(b) $N_{\text{req}}=100$.
(c) $N_{\text{req}}=1000$.
Figure 1. Energy efficiency $\frac{\text{Tokens}}{\text{Joule}}$ of Falcon-RW 7B with different numbers of requests (10, 100, 1000).
Three heatmaps showing the energy efficiency, measured as tokens per joule, of Falcon-RW 7B for 10, 100, and 1000 requests.
Figure 2. LLMs average Energy Efficiency across the entire dataset.
LLMs average Energy Efficiency across the entire dataset.
5.2. Experimental Results
In this subsection we discuss the results of our experimental study, evaluating the energy efficiency of the tested LLMs and the accuracy of our analytical model SweetSpot, using it to predict the peak energy efficiency spots.
(a) XS size models.
(b) S size models.
(c) M size models.
Figure 3. Energy efficiency of tested LLMs considering $n_{\text{in}}=64$ and varying output tokens, per size family (XS, S, M).
@@ 5.2. Experimental Results
5.2.1. Number of Requests Selection
@@
Finally, smaller workloads amplify the relative impact of measurement overhead and necessitate shorter sampling intervals, which can further amplify measurement noise and result in less reliable estimates. For these reasons, in the subsequent experiments we focus on the saturation scenario and fix $N_{\text{req}}$ to 1000. A more detailed analysis of how workload size interacts with batch size is provided in Section 6.2.
5.2.2. Overall LLMs Energy Efficiency
In Figure 2, we report the average energy efficiency measured across the entire dataset for each LLM in our test (see LABEL:tab:llm_set), highlighting the attention mechanism adopted by each model. Among the evaluated models, Llama 3.2 1B achieves the highest overall energy efficiency, with an average score of 55.91. In contrast, Falcon-RW 7B exhibits the lowest efficiency, with an average value of 5.12.
Model size emerges as a primary factor influencing energy efficiency. In general, larger models tend to be less energy efficient than smaller ones. However, parameter count alone does not fully explain the observed differences. The attention mechanism also plays a significant role. In particular, models employing MHA are consistently less energy efficient than comparably sized models adopting GQA. Unlike MHA, where each attention head maintains independent Key–Value (KV) projections, GQA allows multiple query heads to share the same KV projections, which reduces the size of the KV cache and the number of memory accesses required at each decoding step, leading to lower bandwidth pressure and improved energy efficiency. For example, OPT 1.3B (employing MHA) is approximately 42% less efficient than Qwen2 1.5B (GQA 6x), achieving average efficiencies of 20.66 and 49.06, respectively, despite having fewer parameters. A similar pattern is observed when comparing OPT 6.7B and Qwen2 7B. Moreover, higher GQA grouping ratios appear to correlate with improved efficiency. For instance, Qwen 7B (GQA 7x) is only about 18% less efficient than Llama 3.2 3B (GQA 4x), despite having more than twice the number of parameters. This suggests that architectural choices in the attention mechanism can partially mitigate the energy cost typically associated with increased model size.
While the attention mechanism is a critical determinant of energy efficiency, it is not the only architectural factor involved. The number of layers, hidden dimensionality, intermediate projection sizes, and other architectural design choices also substantially impact overall efficiency. A systematic analysis of how these architectural components interact with energy consumption is left for future work.
5.2.3. Observation of Peak Energy Efficiency Spots